### v2

# Latent Inversion with Timestep-aware Sampling for Training-free Non-rigid Editing

Abstract
Text-guided non-rigid editing involves complex edits for input images, such as changing motion or compositions within their surroundings. Since it requires manipulating the input structure, existing methods often struggle with preserving object identity and background, particularly when combined with Stable Diffusion. In this work, we propose a training-free approach for non-rigid editing with Stable Diffusion, aimed at improving the identity preservation quality without compromising editability. Our approach comprises three stages: text optimization, latent inversion, and timestep-aware text injection sampling. Inspired by the recent success of Imagic, we employ their text optimization for smooth editing. Then, we introduce latent inversion to preserve the input image’s identity without additional model fine-tuning. To fully utilize the input reconstruction ability of latent inversion, we suggest timestep-aware text inject sampling. This effectively retains the structure of the input image by injecting the source text prompt in early sampling steps and then transitioning to the target prompt in subsequent sampling steps. This strategic approach seamlessly harmonizes with text optimization, facilitating complex non-rigid edits to the input without losing the original identity. We demonstrate the effectiveness of our method in terms of identity preservation, editability, and aesthetic quality through extensive experiments.
1 Introduction
Diffusion models have become a new standard in image generation and have had a significant impact on text-to-image generation tasks. Text-guided diffusion models take text prompts as input and generate images that align well with the input text prompt. Building upon the success of text-guided diffusion models, there has been active research in text-guided image editing. This encompasses global style transfer Tumanyan et al. (2023); Valevski et al. (2023); Wang et al. (2023b); Zhang et al. (2022) and local editing that require changing object attributes (e.g., color, shape), image translation and adding details Hertz et al. (2022); Couairon et al. (2022); Avrahami et al. (2022); Brack et al. (2023b).
Figure 1: Our method excels detailed non-rigid edits, such as changing the motion or composition of an object on real images directed by text prompts via Stable Diffusion. This figure showcases original images with their corresponding text-guided modification. Our training-free method preserves the original identity and context of the input images while achieving high-level conceptual alterations that seamlessly blend with the images’ inherent characteristics.
Among various local editing tasks, non-rigid editing involves changing the posture or composition of an object while preserving the background and object identity of the original image. It is useful for various applications, such as drawing comic books or videos, where gradual transitions in object actions or compositions over the input image are essential. This task is particularly more challenging than rigid editing, such as global style transfer, image translation, and appearance transformation, which focus on changing color or textures while retaining the image structure. Instead, non-rigid editing requires modifying an object structure partially to align with the text prompt while maintaining its identity.
The challenges of non-rigid editing become more severe when combined with Stable Diffusion Rombach et al. (2022), primarily related to preserving image identity. Prior works in image editing address this problem through various approaches, including injecting attention maps of the source image into the target image generation process Hertz et al. (2022); Tumanyan et al. (2023); Xiao et al. (2023); Cao et al. (2023), model finetuning Ruiz et al. (2023); Valevski et al. (2023); Kawar et al. (2023), and target text optimization Gal et al. (2022); Han et al. (2023); Kawar et al. (2023). However, the attention mechanism often leads to unnatural images as it forcefully combines the structure of the source image with the target, resulting in distorted objects and compositions. Also, model finetuning on an input image can lead to a severe overfitting problem and color distortion, especially when applied to Stable Diffusion.
To this end, our work explores improving identity preservation for non-rigid editing with Stable Diffusion by incorporating a text optimization framework. We focus upon the following key insight: text optimization on an input image has shown promise in enabling smooth changes of an object toward the desired editing concept. However, this hinges on the condition that the model faithfully reproduces the input image and ensures that small changes in text embedding do not significantly alter the output image’s resemblance to the input image Kawar et al. (2023).
Therefore, Imagic Kawar et al. (2023) additionally introduced model finetuning, ensuring the model specialized in reproducing the input image. Nevertheless, as this method causes overfitting and color distortion problems, we make a pivotal change by replacing the model fine-tuning with latent inversion. This change maintains the integrity of the original data distribution, preserving the editability of text optimization. Additionally, we exploit the reconstruction power of latent inversion further by inserting source text embedding exclusively during the initial sampling steps. Then, we transition to target text embedding for the subsequent sampling steps, namely timestep-aware text injection. This concept is inspired by the recent study that reveals the pivotal role of the initial sampling steps in determining the coarse structure and outlines of an image, highly correlated with the identity of input. By selectively injecting source and target prompts considering the role of sampling timesteps, we can explore the smooth transition between identity preservation and editability. This approach allows us to find a sweet spot between these two factors.
Our simple yet effective approach effectively harmonizes with text optimization, achieving complex non-rigid editing without losing the original identity. It does not require additional inputs other than one input image and text prompts. Unlike Imagic, our training-free method offers the advantage of reusing a pre-trained diffusion model regardless of images. Also, our method generates more natural and aesthetically pleasing images than attention-based methods. We compare our edited results both quantitatively and qualitatively with other editing methods, showing superior editing performance on the Stable Diffusion in terms of identity preservation, editability, and aesthetic quality. Moreover, we prove each component of our method is effective through ablation experiments and show the applicability of our method on the different image domain such as Anything-v4.
2 Related Works
2.1 Text-guided Image Manipulation
Text-guided image manipulation involves editing an image based on a given text prompt while preserving the original input image identity. In the realm of this technique, the predominant approach for preserving the original image identity involves the use of an attention mechanism. This mechanism integrates self or cross-attention maps from the input image into the generation process of an edited image Hertz et al. (2022); Tumanyan et al. (2023); Cao et al. (2023); Parmar et al. (2023). Additionally, other methods blend noise maps of the source image and target image, thereby maintaining the background from the source image Couairon et al. (2022); Avrahami et al. (2022); Huang et al. (2023b). Other directions include model finetuning Ruiz et al. (2023); Valevski et al. (2023); Kawar et al. (2023); Li et al. (2023) or text optimization Gal et al. (2022); Han et al. (2023); Kawar et al. (2023), which focus on adapting a model or target text to closely match the input image, even to the extent of overfitting. Furthermore, score distillation has been introduced for image editing, where the input image identity is continuously distilled into the target image during the editing process Hertz et al. (2023). Finally, several approaches Brack et al. (2023a); Tsaban and Passos (2023); Brack et al. (2023b) utilize semantic guidance by adding an edit term in the original classifier-free guidance formula Ho and Salimans (2022).
Figure 2: Ablation. We illustrate the impact of the following components in non-rigid editing: text optimization (Imagic/Ours), model fine-tuning (Imagic), inversion (Ours), and timestep-aware sampling (Ours). The leftmost image is an original input image. The rest of the images are edited according to the target text: “A photo of a white horse jumping” and “A photo of a couple hugging on a beach”. Text optimization minimally affects identity preservation. Model fine-tuning from Imagic enhances image resemblance to the input, often at the expense of color fidelity. Conversely, our latent inversion stabilizes the overfit tradeoff and avoids color distortion. To further improve identity preservation, we incorporate source text embedding during the initial sampling timesteps.
2.2 Diffusion Inversion
For real image editing, the concept of inversion has been widely used to uncover the latent representation that faithfully reconstructs the input real image. Inversion methods in the diffusion model have been evolved in order to improve both reconstruction and editing ability. DDIM inversion Song et al. (2020) excels at reconstructing the original image for unconditional image generation scenarios. However, it faces challenges in conditional cases where the use of classifier-free guidance (CFG) Ho and Salimans (2022) magnifies the errors, resulting in decreased editability and reconstruction quality. This drawback of DDIM gave rise to active research in inversion specialized for conditional image generation. Null-text inversion Mokady et al. (2023) improves the inversion quality by optimizing null text embedding $\emptyset$ at each diffusion timestep. On the other hand, Direct Inversion Ju et al. (2023) takes a different approach by disentangling source and target diffusion branches, eliminating optimization. This strategy boosts content preservation quality and edit fidelity within a short time. EDICT Wallace et al. (2023) offers a mathematically exact inversion solution through an alternative inversion process that involves a pair of coupled noise vectors. Edit-friendly inversion Huberman-Spiegelglas et al. (2023) strongly imprints the input image onto latent noise maps by making them correlated across timesteps, enabling fast and effective editing without the need for optimization. More explanations on the diffusion process, classifier-free guidance and inversion methods that we utilized are in Appendix.
3 Preliminary
Stable Diffusion Model. Diffusion models aim to estimate a data distribution via a latent variable model. They perform a forward process that gradually adds a Gaussian noise on a clean image at each timestep $t\in[1,...,T]$ until it reaches a noise distribution. Conversely, the reverse process inversely denoises the noise distribution in order to generate a final clean image. The Stable Diffusion model Rombach et al. (2022), an open-source model, conducts these processes on the latent space. Specifically, given an image $\bm{x}\in\mathbb{R}^{H\times W\times 3}$, it encodes $\bm{x}$ into a latent representation $\bm{z}_{0}=\mathcal{E}(\bm{x})$ using a VQ-GAN Esser et al. (2021) encoder $\mathcal{E}$. The model is trained on this latent space, applying the forward process to $\bm{z}_{0}$. Likewise, $\bm{z}_{0}$ is obtained from the latent variable $\bm{z}_{T}$ using the reverse process. Finally, the generated image is computed by decoding $\bm{z}_{0}$ back to the image space ($\bm{x}=\mathcal{D}(\bm{z}_{0})$). Also, an important function of Stable Diffusion is its capability to facilitate text-guided image generation. This is achieved by inputting a text embedding $\tau_{\phi}(\bm{y})$, which is a text prompt $\bm{y}$ processed by CLIP Radford et al. (2021) text encoder $\tau_{\phi}$. The text-guided conditional model is trained using the following objective:
$\small\mathcal{L}_{\text{cond}}=\mathbb{E}_{t,\bm{z}_{0},\bm{\epsilon}}\left[||\bm{\epsilon}-f_{\theta}(\bm{z}_{t},t,\tau_{\phi}(\bm{y}))||^{2}\right],$ (1)
where $f_{\theta}$ is a neural network with parameter $\theta$, $t\sim\texttt{Uniform}[1,T]$, $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$, and $\bm{z}_{t}$ is a noise added version of $\bm{z}_{0}$ using the forward process. Our goal is to improve non-rigid editing with Stable Diffusion, aiming to enhance the accessibility of a publicly available model. Detailed explanations of the forward and reverse process of diffusion models are available in the Appendix.
Challenges in Non-rigid Editing. Compared to rigid editing, such as global style transfer, image translation, and appearance transformation, non-rigid editing stands out as highly challenging. Specifically, rigid editing handles relatively simple color adjustments while maintaining the image structure intact, which model filters can easily capture. Conversely, non-rigid editing involves understanding and partially modifying an underlying structure of the original object while maintaining its identity. For that, the model should faithfully reconstruct the original input while selectively editing the structural components that align with the target prompt. Since faithful reconstruction has a trade-off relationship with diverse editing in most text-guided image generation models, non-rigid editing often lacks identity preservation. Therefore, improving the identity preservation ability of the model is the main issue in enhancing the non-rigid editing performance.
Figure 3: Overview. Our method sequentially performs (A) text optimization, (B) latent inversion, and (C) timestep-aware injection sampling. (A) The target text is encoded into an optimized embedding to match the input image. (B) Null-text inversion is employed for additional identity preservation. (C) The source text embedding $\bm{e}_{src}$ instead of an interpolated embedding $\bm{e}_{int}$ is injected during the initial sampling steps to anchor the image’s global structure more effectively.
Imagic for Non-rigid Editing. Imagic Kawar et al. (2023), a pioneering method for non-rigid editing, introduced a two-step process: text optimization and model fine-tuning. Firstly, they optimize the target text embedding to best match the input image. Specifically, the target text is encoded to a text embedding $\bm{e}_{tgt}\in\mathbb{R}^{N\times d}$ via a text encoder Raffel et al. (2020), where $N$ is the number of tokens in the target text and $d$ is the token embedding dimension. Following this, they freeze parameters of the pre-trained diffusion model $f_{\theta}$ and optimize the target text embedding $\bm{e}_{tgt}$ using the diffusion model objective:
$\small\bm{e}_{opt}=\arg\min\mathbb{E}_{t,\bm{z}_{0},\bm{\epsilon}}\left[\left\|\bm{\epsilon}-f_{\theta}\left(\bm{z}_{t},t,\bm{e}_{tgt})\right)\right\|^{2}\right].$ (2)
Though the target text embedding is optimized to be matched with the input image, $\bm{e}_{opt}$ often does not show a perfect representation of the input image. Additional optimization does not improve identity preservation but causes a runtime overhead problem (Text optim step 200, Text optim step 1000 in Figure 2). Therefore, Imagic uses model fine-tuning as a firm identity preservation strategy to make the model overfitted to the input image. It freezes the optimized embedding and then fine-tunes the model parameters $\theta$ using the same loss function in Equation 2.
Finally, during the sampling phase, Imagic linearly interpolates the original target text embedding with the optimized one using Equation 3 and pass this through the model to generate an edited image. Here, it can control the degree of editing by varying an interpolation ratio $\alpha$. This aims to establish a smooth transition of text embedding toward the desired editing concept while maintaining the original identity.
$\bm{e}_{int}=\alpha\cdot\bm{e}_{tgt}+(1-\alpha)\cdot\bm{e}_{opt}$ (3)
It is worth noting that Imagic is particularly effective when applied to Imagen Saharia et al. (2022), a large-scale generative model only accessible for business. It successfully applies the editing concept while faithfully maintaining the original identity and background of the original image. However, the same recipe encounters limitations when applied to Stable Diffusion, as mentioned in previous works Han et al. (2023); Brack et al. (2023b). Our experiments also affirm the unsatisfactory editing performance of Imagic on Stable Diffusion (Sec 5), with a noticeable deficiency in identity preservation.
4 Method
In this section, we identify the problem of Imagic combined with Stable Diffusion and introduce our key contributions to enhancing non-rigid editing performance. An overview of our method is demonstrated in Figure 3.
We follow the text optimization from Imagic to preserve the original identity and achieve smooth non-rigid editing (See Figure 3(A)). As text optimization alone is insufficient for a perfect match with the input image, Imagic solves this problem via model finetuning. However, in the Stable Diffusion setting, we found that the model finetuning presents several significant drawbacks that adversely affect editing quality. Firstly, it triggers a severe tradeoff between (1) overfitting, where the model overly memorizes the original image and thus cannot faithfully incorporate the text prompt well, or (2) the lack of identity preservation, where the edited image aligns with the target text but it fails to preserve the original image’s identity. Figure 2 provides visual evidence that the edited image closely resembles the input image as we finetune the model longer. Conversely, minimal finetuning produces an arbitrary image that bears little resemblance to the input image, indicating the lack of identity preservation (Text optim step 200 + Finetune step 100, +Finetune step 400 in Figure 2). Model finetuning thus hardly satisfies both identity preservation and editability simultaneously. (More ablation experiments on finetuning steps in the Appendix.) Even worse, it causes color distortion, disrupting the data distribution. This is because model training aims to reach the data distribution of the training dataset, which in this context consists of only a single input image.
Instead of model fine-tuning, we utilize latent inversion to maintain the original identity (See Figure 3(B)). The inversion process is defined by Equation 4:
$\small\bm{z}^{*}_{t+1}=$ $\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}\bm{z}^{*}_{t}+$
$\sqrt{\alpha_{t+1}}\left(\sqrt{\frac{1}{\alpha_{t+1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\right)f_{\theta}(\bm{z}^{*}_{t},t,\bm{e}_{src}),$ (4)
where $\alpha_{t}$ is a noise schedule at timestep $t$ and $\bm{e}_{src}$ is a source text embedding. As the inversion produces a latent that faithfully reconstructs the original image, the use of inverted latent for sampling can effectively maintain the original identity with text-guided editing. Moreover, the inversion method no longer suffers from (1) the tradeoff between overfitting and editability and (2) color distortion.
While employing the inverted noise helps maintain the overall color and context of the original image, introducing a new text prompt different from the source text can cause deviations in image configuration, such as the size and position of the object, compared to the original image. To address this issue, we leverage a well-known observation that diffusion models generate global structures at the initial timesteps and fine details toward the end Choi et al. (2022). Specifically, we devise an additional regularization depending on sampling timesteps (See Figure 3(C)). During the initial steps ($n\rightarrow T$), we incorporate the source text embedding and then transition to interpolated embeddings for the subsequent timesteps:
[…]
This strategy helps to closely adhere to the global structure of the original image, a key component of the image identity, during the early stage of the editing process. Since then, the editing concept has been applied to the subsequent sampling steps, effectively incorporating editing concepts without altering the image identity. We show its substantial impact on preserving the identity through ablation studies in Section 5.4.
Figure 4: Qualitative results. We compare our editing results to five recent editing methods. While other methods suffer from either overfitting the input image or lacking identity preservation, our method shows the best structure preservation and edit fidelity.
5 Experiments
5.1 Implementation Details
Dataset. We employ the Textual Editing Benchmark (TEdBench), introduced by Imagic Kawar et al. (2023). It comprises input images with target texts describing desired complex non-rigid editing.
Baselines. For a comprehensive comparison, we evaluate our method against the following competitive techniques: Null-text inversion Mokady et al. (2023), Edit Friendly Huberman-Spiegelglas et al. (2023), Imagic Kawar et al. (2023), DDS Hertz et al. (2023), and Masactrl Cao et al. (2023). Notably, Imagic and Masactrl are specialized for the non-rigid editing task. Null-text inversion and Edit-Friendly represent prominent methods designed for natural image editing via inversion. DDS proposes a new paradigm of image editing using score distillation. In our evaluations, we run the released codes of all baseline models using pre-trained Stable Diffusion v1.4 and anime-style model Anything-V4 with publicly available checkpoints. Unfortunately, we could not include other non-rigid editing methods, such as LayerDiffusion Li et al. (2023) and KVInversion Huang et al. (2023a), as their codes are not released as of our submission date. For sampling, we use a conventional setting of DDIM with 50 timesteps  Cao et al. (2023); Huang et al. (2023a); Li et al. (2023) for all methods. For our method using DDIM inversion Song et al. (2020), we used a guidance scale of 1 for both inversion and sampling stages, as it is known that using a larger guidance scale leads to suboptimal results. For our method using Null-text inversion Mokady et al. (2023), we used a guidance scale of 7.5 for both inversion and sampling following previous works Cao et al. (2023); Huang et al. (2023a).
Evaluation Metrics. For measuring target text alignment and input image identity preservation quality of edited images, we utilize CLIP-T score Radford et al. (2021) and LPIPS score Zhang et al. (2018), following conventions Kawar et al. (2023); Mokady et al. (2023). CLIP score concerning the target text measures the cosine similarity between the embeddings of the target text prompt and the edited image. LPIPS score measures the difference between the original input image and the edited image in the VGG Simonyan and Zisserman (2014) feature space. However, these two metrics are incapable of evaluating the aesthetic quality of the edited image; they cannot penalize the distorted or unnatural outputs. To mimic the standard of human visual inspection, we additionally employ an Aesthetic score Schuhmann (2022), which measures how much people like the image on average. Few MLP layers are added on the CLIP ViT-L/14, and only the MLP layers are finetuned to predict the aesthetic score using MSE or MAE loss. AVA Murray et al. (2012) and Simulacra Pressman et al. (2022) are used for training the predictor, which are datasets that collect human ratings of 1 to 10 on real and generated images each. This metric is used to clean the LAION-5B Schuhmann et al. (2022) dataset.
5.2 Comparison with Other Works
Results with Stable Diffusion. We compare the editing results of our method to other baselines using a pre-trained Stable Diffusion v1.4 in Figure 4. We show our results using Null-text inversion as it provides the best editing performance (Figure 7). For a fair comparison, we also applied Null-text inversion in Masactrl, which originally used DDIM inversion. It results in improved identity preservation and editability of Masactrl. We tried various hyperparameters of each method and chose the best result for each.
Firstly, Null-text inversion alone for editing, in which we just input an inverted image and a target text through the model, shows low identity preservation. It does not retain the original structure of the input image and simply pursues generating an image that aligns well with the text prompt. This is expected because there is no other constraint for identity preservation during the sampling process. Meanwhile, Imagic has a serious tradeoff between overfitting and editability. Model finetuning especially causes a severe overfitting problem, which distorts the data distribution, whereas a random noise without finetuning during the sampling process ruins the structure preservation. Despite the extensive hyperparameter search, it is hard to find a sweet spot between these two factors (See Appendix). Besides, DDS also suffers from overfitting to the input image, especially when handling non-rigid editing. This shows that score distillation has trouble changing geometry. On the other hand, Edit-Friendly and Masactrl attempt to apply an editing concept while maintaining the original identity by improving latent noise space and utilizing attention maps, respectively. When an image has no complicated background and contains a single object, such as a Spiderman image in the third row of Figure 4, their editing quality is sufficiently high. However, they show unnatural edited results in other cases. In the first row of Figure 4, a bulldog opens its mouth but does not smile in Edit-Friendly, and the bulldog shows a grotesque mouth in Masactrl. In the second row, a horse jumps in both methods, but the hind legs of the horse remain similar to the input image, resulting in an unnatural jumping appearance. Overall, our method clearly shows the editing concept while faithfully preserving the original identity, like a bulldog with a big smile and a horse jumping dynamically. These visual trends and the superiority of our method consistently hold in other results. More qualitative results are demonstrated in Appendix.
Results with Anything-V4. We also perform our method on a pretrained anime-style model, Anything-V4. We generated an image using the model and then performed editing just like the real image editing setting. Figure 5 shows various editing results using diverse text prompts. We can change the posture, facial expression, and viewpoint smoothly while maintaining the original structure of the input image. This proves that our method generalizes well regardless of the image domain. More qualitative results are shown in Appendix.
Figure 5: Anything-v4. This figure showcases the adaptability of our model on an Anything-V4 anime-style model. Our method effectively edits the input image’s posture and viewpoint with various text prompts, smoothly preserving the original structure.
5.3 Quantitative Evaluation
We conduct a quantitative evaluation using the following non-rigid editing models: Imagic, Masactrl, and the proposed method. The experiment is carried out using the pre-trained Stable Diffusion v1.4. To ensure a fair comparison, we use Null-text inversion for both ours and Masactrl. We utilized the TedBench dataset to generate 86 edited images, producing five variants for each edited image with different seeds, resulting in a total of 430 images. In Figure 6, we present CLIP and Aesthetic scores of each method at different LPIPS scores, where LPIPS changes depending on hyperparameters. Herein, a higher CLIP score refers to better alignment with target text while keeping the LPIPS score fixed. Imagic consistently yields the lowest CLIP score across all settings, indicating that its editability stays far behind ours and Masactrl. Our method exhibits the highest CLIP score in most settings. It implies that our method effectively preserves the input image’s identity while accurately conveying the editing concept compared to other methods. Regarding the Aesthetic score, our method achieves the highest score in all the settings. It highlights that our method has both superior editability and aesthetic appeal compared to other methods. This success can be attributed to our inversion and timestep-aware text injection sampling, which preserve the original identity without disrupting the data distribution. This enables text optimization to smoothly transition towards the desired editing concept while maintaining the background and object identity faithfully. Imagic tends to overfit the input image, maintaining the naturalness of the image structure, whereas color gets distorted. Consequently, its aesthetic score lags behind ours, but not the least. Finally, Masactrl presents the lowest Aesthetic score, consistent with its tendency to generate unnatural images with distorted object appearances and poses. This can be attributed to the difference in object appearance and poses between the source and target images, each generated using the original source and target text, respectively. Example images are available in the Appendix. As the appearance gap between the target image and the source image intensifies, inserting the attention maps of the source image into the target image sampling process often results in abnormal objects with unnatural appearances and poses. Though there is a point where Masactrl has a higher CLIP score compared to ours, its low Aesthetic score shows that it has difficulty generating an edited image in which the editing concept seamlessly blends with the object’s intrinsic characteristics.
Figure 6: Quantitative results. The graph represents (A) CLIP scores ($\uparrow$), indicating alignment with target text, and (B) Aesthetic scores ($\uparrow$), under various image quality levels as LPIPS scores change. Our method outperforms Imagic and Masactrl in both CLIP and Aesthetic scores in most settings, demonstrating superior editability and aesthetic quality. This highlights the effectiveness of our method, latent inversion with timestep-aware text injection sampling, in preserving the original identity and data distribution, leading to more natural and concept-aligned edits.
5.4 Ablation Experiments
We show the effectiveness of each component of our method through ablation experiments on inversion type and source text injection ratio.
Inversion Type. Figure 7 demonstrates the impact of different inversion methods: DDIM Song et al. (2020), Null-text inversion Mokady et al. (2023) with (NTI-src) and without source prompt (NTI-w/o src). Each image is edited with the target text “A photo of a jumping dog.” When employing DDIM inversion for our method, it is highly close to the input image but has low editability. This again proves that DDIM inversion is suboptimal for real image editing applications. On the other hand, Null-text inversion is known to improve inversion and editing quality, particularly when using classifier-free guidance with a guidance scale greater than 1, as compared to DDIM inversion. Interestingly, when using the source prompt, the preservation quality sometimes significantly deteriorates while focusing on editing (NTI - src). Conversely, when we set a source text prompt equal to null text, the edited images show significant improvement in structural preservation (NTI - w/o src). Therefore, when the reconstruction performance of inversion falls behind, employing the null text for the source prompt can enhance the overall reconstruction ability. Overall, Null-text inversion demonstrates superior performance in editing quality to DDIM.
Figure 7: Ablation on inversion type. We demonstrate our editing results with three inversion types: DDIM, Null-text inversion with source (NTI-src), and without source embedding (NTI-w/o src). The results confirm that DDIM is less favorable to the editing task than NTI, where setting null text instead of source text often produces more plausible results.
Source Text Injection Ratio. Our timestep-aware sampling strategy injects the source text embedding at the beginning of the sampling process instead of an interpolated embedding to improve identity preservation. Figure 2 includes the ablation study with different source text input ratios. Given that we use DDIM sampling with 50 steps, inserting the source text for the initial 10 steps corresponds to a 20% source text input ratio. We present results under three different scenarios: a ratio of 0%, 10%, and 20%, depicted as Text optim step 200 + Inversion, + Src ratio 10%, and + Src ratio 20% in Figure 2, respectively. Notably, as the source text input ratio increases, the edited images increasingly retain the original structure of the input image. We empirically observe that inserting the source text about 5 to 20 steps, which corresponds to a ratio of 10% to 40%, leads to the most favorable editing results depending on the degree of desired structural changes.
6 Conclusion
Our work has advanced the capabilities of non-rigid editing with Stable Diffusion, particularly focusing on identity preservation. Our method comprises text optimization, latent inversion, and timestep-aware text injection sampling. Based on the analysis of the previous study, we chose text optimization as the baseline as it allows a smooth transition toward a desirable editing concept. To overcome the limitations of model fine-tuning from Imagic, we suggest utilizing the reconstruction power of latent inversion. To further improve the identity preservation quality, we devise timestep-aware text injection sampling that selectively applies source and target prompts depending on sampling timesteps for improved identity preservation and editability. Our method presents superior performance in both qualitative and quantitative comparisons with other editing methods on Stable Diffusion. Our method does not trigger overfitting tradeoffs and color distortion, unlike model finetuning in Imagic. Also, we can provide natural and aesthetically pleasing images compared to attention-based methods. Finally, we can extend the usage of our method for other image domains like Anything-V4.
Appendix A Diffusion Models
A.1 Denoising Diffusion Probabilistic Models
Diffusion models are latent variable models that generate an image $\bm{z}_{0}\sim q(\bm{z}_{0})$ by iteratively denoising a noise distribution. They perform forward and reverse processes defined as a Markov chain. A forward process samples $\bm{z}_{t}$ at timestep $t\in[1,...,T]$ by gradually adding a Gaussian noise $\beta_{t}\in(0,1)$ starting from $\bm{z}_{0}$:
$q(\bm{z}_{1},...,\bm{z}_{T}|\bm{z}_{0})$ $\coloneqq\prod_{t=1}^{T}q(\bm{z}_{t}|\bm{z}_{t-1}),$
$q(\bm{z}_{t}|\bm{z}_{t-1})$ $\coloneqq\mathcal{N}(\bm{z}_{t};\sqrt{1-\beta_{t}}\bm{z}_{t-1},\beta_{t}\mathbf{I}).$ (6)
We can obtain $\bm{z}_{t}$ directly from $\bm{z}_{0}$ as follows:
$\small\bm{z}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},$
where $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, $\alpha_{t}=1-\beta_{t}$, and $\bar{\alpha}_{t}=\prod_{i=0}^{t}\alpha_{i}$. The latent $\bm{z}_{T}$ becomes nearly an isotropic Gaussian distribution under the assumption of large $T$ and a well-defined $\beta_{t}$. Therefore, to generate an image, we can sample $\bm{z}_{T}\sim\mathcal{N}(0,\mathbf{I})$ and obtain $\bm{z}_{0}$ through a exact reverse process $q(\bm{z}_{t-1}|\bm{z}_{t})$. However, as $q(\bm{z}_{t-1}|\bm{z}_{t})$ relies on the entire data distribution, we utilize a neural network with parameter $\theta$ to approximate it:
$p(\bm{z}_{0},...,\bm{z}_{T})$ $\coloneqq p(\bm{z}_{T})\prod_{t=1}^{T}p(\bm{z}_{t-1}|\bm{z}_{t}),$
$p_{\theta}(\bm{z}_{t-1}|\bm{z}_{t})$ $\coloneqq\mathcal{N}(\bm{z}_{t-1};\mu_{\theta}(\bm{z}_{t},t),\Sigma_{\theta}(\bm{z}_{t},t)).$
The reverse process becomes tractable when conditioned on $\bm{z}_{0}$ using Bayes theorem:
$q(\bm{z}_{t-1}|\bm{z}_{t},\bm{z}_{0})$ $=\mathcal{N}(\bm{z}_{t-1};\tilde{\mu}(\bm{z}_{t},\bm{z}_{0}),\tilde{\beta}_{t}\mathbf{I}),$
$\tilde{\mu}_{t}(\bm{z}_{t},\bm{z}_{0})$ $\coloneqq\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\bm{z}_{0}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\bm{z}_{t},$
$\tilde{\beta}_{t}$ $\coloneqq\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}.$
Finally, the neural network is trained using the variational lower bound (VLB):
$L_{\text{vlb}}$ $\coloneqq L_{0}+L_{1}+...+L_{T-1}+L_{T},$
$L_{0}$ $\coloneqq-\log p_{\theta}(\bm{z}_{0}|\bm{z}_{1}),$
$L_{t-1}$ $\coloneqq D_{KL}(q(\bm{z}_{t-1}|\bm{z}_{t},\bm{z}_{0})\;||\;p_{\theta}(\bm{z}_{t-1}|\bm{z}_{t})),$
$L_{T}$ $\coloneqq D_{KL}(q(\bm{z}_{T}|\bm{z}_{0})\;||\;p(\bm{z}_{T})).$
It is available for the network to predict either $\bm{z}_{0}$ or $\bm{\epsilon}$, proving that the latter provides higher sample quality empirically. Therefore, the VLB loss gets simplified into $\bm{\epsilon}$-prediction loss:
$\small L_{\text{simple}}=E_{t,\bm{z}_{0},\bm{\epsilon}}\left[||\bm{\epsilon}-f_{\theta}(\bm{z}_{t},t)||^{2}\right],$
where $t\sim\texttt{Uniform}[1,T]$.
After the training, we can generate an image $\bm{z}_{0}$ through a reverse process which iteratively denoises the latent starting from the latent $\bm{z}_{T}\sim\mathcal{N}(0,\mathbf{I})$. Specifically, we obtain $\bm{z}_{t-1}$ from $\bm{z}_{t}$ by inserting a predicted epsilon, $f_{\theta}(\bm{z}_{t},t)$ into the Equation (7):
$\small\bm{z}_{t-1}=\frac{1}{\sqrt{1-\beta_{t}}}\left(\bm{z}_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{\alpha}_{t}}}f_{\theta}(\bm{z}_{t},t)\right)+\sigma_{t}\bm{\epsilon},$ (7)
where $\sigma_{t}^{2}=\beta_{t}$.
Figure 8: Ablation on Model Finetuning in Stable Diffusion. Few finetuning steps yields higher editability with low identity preservation quality, whereas longer finetuning triggers overfitting and color distortion problem. It is difficult to find a sweet spot of the tradeoff between identity preservation and editability.
A.2 Denoising Diffusion Implicit Models
DDPMs have shown remarkable performance in image generation. However, as the sampling process is based on a Markov chain, it takes a long time to produce one image. DDIM addresses this problem by encompassing DDPM into a class of non-Markovian process while maintaining the same training objective. Among available inference distributions indexed by a real vector $\sigma\in\mathbb{R}_{\geq 0}^{T}$, we can define
$q_{\sigma}(\bm{z}_{1:T}|\bm{z}_{0}):=q_{\sigma}(\bm{z}_{T}|\bm{z}_{0})\prod_{t=2}^{T}q_{\sigma}(\bm{z}_{t-1}|\bm{z}_{t},\bm{z}_{0}),$
$q$ (${}_{\sigma}(\bm{z}_{t-1}|\bm{z}_{t},\bm{z}_{0})$
$=\mathcal{N}\left(\sqrt{\alpha_{t-1}}\bm{z}_{0}+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\cdot{\frac{\bm{z}_{t}-\sqrt{\alpha_{t}}\bm{z}_{0}}{\sqrt{1-\alpha_{t}}}},\sigma_{t}^{2}\mathbf{I}\right),$ (8)
where $q_{\sigma}(\bm{z}_{T}|\bm{z}_{0})=\mathcal{N}(\sqrt{\alpha_{T}}\bm{z}_{0},(1-\alpha_{T})\mathbf{I})$.
The Equation (8) is formed to coincide with the original marginals of DDPMs, $q_{\sigma}(\bm{z}_{t}|\bm{z}_{0})=\mathcal{N}(\sqrt{\alpha_{t}}\bm{z}_{0},(1-\alpha_{t})\mathbf{I})$ for all $t$. Then, we can calculate the forward process using Bayes theorem:
$q_{\sigma}(\bm{z}_{t}|\bm{z}_{t-1},\bm{z}_{0})=\frac{q_{\sigma}(\bm{z}_{t-1}|\bm{z}_{t},\bm{z}_{0})q_{\sigma}(\bm{z}_{t}|\bm{z}_{0})}{q_{\sigma}(\bm{z}_{t-1}|\bm{z}_{0})}.$
Unlike the forward process in Equation (6), $\bm{z}_{t}$ is now conditioned on both $\bm{z}_{t-1}$ and $\bm{z}_{0}$, resulting in non-Markovian forward process.
Likewise, the reverse process is reformulated as:
$\bm{z}_{t-1}$ $=\sqrt{\alpha_{t-1}}\underbrace{\left(\frac{\bm{z}_{t}-\sqrt{1-\alpha_{t}}f_{\theta}(\bm{z}_{t},t)}{\sqrt{\alpha_{t}}}\right)}_{\text{`` predicted }\bm{z}_{0}\text{''}}$
$+\underbrace{\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot f_{\theta}(\bm{z}_{t},t)}_{\text{``direction pointing to }\bm{z}_{t}\text{''}}+\underbrace{\sigma_{t}\bm{\epsilon}}_{\text{random noise}}.$ (9)
When $\sigma_{t}=\sqrt{\left(1-\alpha_{t-1}\right)/\left(1-\alpha_{t}\right)}\sqrt{1-\alpha_{t}/\alpha_{t-1}}$ for all $t$, the forward process becomes Markovian, representing DDPM. When $\sigma_{t}=0$, the forward process becomes deterministic, enabling accelerated sampling.
A.3 Classifier-free Guidance
The main goal of text-guided generative models is to magnify the impact of an input text. Classifier-free guidance (CFG) Ho and Salimans [2022] enables this by incorporating the use of null text embedding $\emptyset=\tau_{\phi}(``")$, the null text processed by CLIP Radford et al. [2021] text encoder $\tau_{\phi}$. Specifically, an epsilon predicted from null text embedding $\emptyset$ is extrapolated with the one predicted from an input text embedding $\bm{e}$ via a guidance scale $w$:
$\tilde{f}_{\theta}(\bm{z}_{t},t,\bm{e},\emptyset)=w\cdot f_{\theta}(\bm{z}_{t},t,\bm{e})+(1-w)\cdot f_{\theta}(\bm{z}_{t},t,\emptyset)$ (10)
The guidance scale enables a tradeoff between sample fidelity and diversity. Images generated with a higher guidance scale align more faithfully with the input text.
Appendix B Diffusion Inversion
B.1 DDIM Inversion
The inversion (forward) process of DDIM is represented by inverting the reverse process:
$\bm{z}_{t+1}=$ $\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}\bm{z}_{t}+$
$\sqrt{\alpha_{t+1}}\left(\sqrt{\frac{1}{\alpha_{t+1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\right)f_{\theta}(\bm{z}_{t},t,\bm{e}).$
DDIM inversion Song et al. [2020] works well for unconditional image generation in which CFG Ho and Salimans [2022] is not used ($w=1$), although a slight error occurs each step. However, it does not work well in conditional cases as CFG magnifies the errors, decreasing editability and reconstruction quality. Since CFG is useful for conditional image generation, inversion methods to faithfully reconstruct the images when using CFG have been actively studied.
B.2 Null-text Inversion
Null-text inversion Mokady et al. [2023] improves the inversion quality using CFG by optimizing null text embedding $\emptyset$ at each diffusion timestep. Since DDIM inversion Song et al. [2020] with guidance scale $w=1$ offers a promising guideline for inversion, they calculate $(\bm{z}^{*}_{t})_{t\in\{1,\cdots,T\}}$ from $\bm{z}_{0}$ via DDIM inversion Song et al. [2020] and set these as pivot trajectories. They initialize $\bar{\bm{z}}_{T}=\bm{z}^{*}_{T}$ and $\varnothing_{T}=\varnothing$. Then, they calculate $\bm{z}_{t-1}$ from $\bar{\bm{z}}_{t}$ using DDIM (9), CFG (10) and null-text embedding $\varnothing_{t}$ at each diffusion step $t$:
$\bm{z}_{t-1}$ $(\bar{\bm{z}}_{t},t,\bm{e},\varnothing_{t})=\sqrt{\frac{\alpha_{t-1}}{\alpha_{t}}}\bar{\bm{z}}_{t}+$
$\sqrt{\alpha_{t-1}}\left(\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\right)\tilde{f}_{\theta}(\bar{\bm{z}}_{t},t,\bm{e},\varnothing_{t}).$
They optimize $\varnothing_{t}$ via a MSE loss between $\bm{z}_{t-1}(\bar{\bm{z}}_{t},t,\bm{e},\varnothing_{t})$ and $\bm{z}^{*}_{t-1}$:
$\min_{\varnothing_{t}}\|\bm{z}_{t-1}(\bar{\bm{z}}_{t},t,\bm{e},\varnothing_{t})-\bm{z}^{*}_{t-1}\|^{2}_{2},$
They update $\bar{\bm{z}}_{t-1}=\bm{z}_{t-1}(\bar{\bm{z}}_{t},t,\bm{e},\varnothing_{t})$ after the optimization of each timestep. For our experiments, we follow the default hyperparameter of Null-text inversion, optimizing 10 iterations for each timestep.
Figure 9: Ablation on Source Text Injection Ratio. Low injection ratio leads to higher editability while high ratio more preserves the original identity. The optimal ratio depends on the desired structural changes, usually ranging from 10% to 40%.
Figure 10: Failure Cases of Attention Mechanism. Attention mechanisms sometimes result in unnatural images due to the gap between source image and target image.
Appendix C Limitation of Model Finetuning
We show the impact of model finetuning in Stable Diffusion Rombach et al. [2022] as the finetuning step increases by 50 in Figure 8. Each row represents the edited results according to the target text: “A photo of a white horse jumping” and “A photo of a couple hugging on a beach.” In the early steps, the images align well with the target text but lack object identity and background preservation (Step 50). As we finetune more to maintain the original identity, the edited images get strongly overfitted to the input image, no longer revealing the desired editing concept. Furthermore, the color gets distorted, resulting in images with high-frequency artifacts (Step 300, Step 350, Step 400). The output images at the middle steps do not satisfy both identity preservation and editability (Step 150, Step 200, Step 250). It indicates that it is hard to find a sweet spot of a tradeoff between identity maintenance and editability of model finetuning. Occasionally, it even takes a long time to get overfitted to the input image, as seen in the second row of Figure 8. The position of the couple and rocks mismatch with the input image while the color distortion has already begun (Step 400).
Appendix D Limitation of Attention Mechanism
The attention mechanism often produces unnatural images with distorted object appearances and poses. It is due to the gap between object appearances and poses of source image and target image, which are generated from source text and target text. Figure 10 showcases the corresponding examples. The first two columns are the source and target image, respectively, and the last column is the edited result of Masactrl Cao et al. [2023]. The source and target image in the first row differ in terms of the background, size, and color of the cats. Those in the second row also have a completely different appearance of streets. Nevertheless, the attention mechanism injects the attention map of the source image forcefully into the target image during the generation process of the edited image. This results in unnatural, aesthetically unsatisfactory images, as shown in the last column.
Figure 11: Qualitative Comparison. Our method has better identity preservation quality and editability compared to other editing methods.
Figure 12: Limitations. Failure cases on (A) interaction between multiple objects and (B) compositions.
Appendix E More Experiments
E.1 Ablation on Source Text Injection Ratio
We inject source text embedding at the beginning of the sampling process, and then transition to the interpolated embedding later. Figure 9 visualizes edited results under various injection ratios. As we utilize DDIM Song et al. [2020] sampling step 50, injection ratio 10% means injecting the source embedding for the first five steps. With the low ratio, the editing concept is represented in the image, but the identity preservation quality is unsatisfactory. Due to the use of inversion, the images tend to maintain the identity of the original image to some extent, but there is still some space for improvement (Src ratio 0%). For example, the background of the cat image, the size and position of the couple, and the color of the door mismatch with the input image. As the ratio increases, preservation quality is enhanced, resulting in overfitting when it goes too far (Src ratio 50%). The optimal injection ratio varies depending on the degree of desired structural changes. When a significant change is required, like the first row of Figure 9, 10% of injection is adequate. As the structural change decreases, the optimal ratio increases as an increasing effect of source embedding is needed for identity preservation. Roughly speaking, 40% injection is required for the simple editing case, such as adding the number on the input image. Overall, the injection ratio of 10% to 40% is effective for satisfying both identity preservation and editability.
E.2 Qualitative Comparison
We show more qualitative comparison results in Figure 11. We compare our method with the recent five editing methods. We tried various hyperparameters of each method and chose the best result for each. Our method satisfies both identity preservation quality and editability compared to other methods. We provide more qualitative results using the pre-trained Stable Diffusion and Anything-v4 models in Figure 13 and Figure 14. Our method can perform non-rigid editing encompassing various changes of motion or composition of objects.
Appendix F Limitations
Although our method can effectively perform non-rigid editing, we entail limitations of compositional image generation of Stable Diffusion Rombach et al. [2022]. We have some difficulties when dealing with compositions or interactions between multiple objects. Figure 12 visualizes several failure cases. In (A), the text prompts for making two parrots look at each other and a goat jumping over a cat were not properly reflected. Also, compositions of objects represented with prepositions are non-trivial, as evident in (B). That is, the intention was to generate a cup next to a vase, but the model transforms the vase into the cup. Similarly, a leaf was intended to be on top of a book, but it appears under the book. These limitations arise from the mismatch between attention regions and each object Wang et al. [2023a], incapable of exactly delineating the boundary of each attention region. Consequently, artifacts become evident, especially for images comprising multiple objects. The same problem also appears in other editing methods that we used for comparison with ours.
Also, there is potential for improving identity preservation. Our method maintains the overall identity of the input image, but sometimes, it may miss high-frequency details. Finally, our method introduces a runtime overhead for editing, primarily due to the need for text optimization as well as null-text optimization in Null-text inversion. Herein, we specifically use 200 text optimization steps. When employing source text for inversion, our method takes about 1.5 minutes. Conversely, setting the source text as null text takes about 1 minute. We can reduce the runtime by using other inversion methods that do not require optimization and, therefore, directly perform the inversion.
Figure 13: Qualitative results on Stable Diffusion.
Figure 14: Qualitative results on Anything-v4.