### v3

# Daejeon, Republic of Korea BMVC Author Guidelines

Abstract
Text-guided non-rigid editing involves complex edits for input images, such as changing motion or compositions within their surroundings. Since it requires manipulating the input structure, existing methods often struggle with preserving object identity and background, particularly when combined with Stable Diffusion. In this work, we propose a training-free approach for non-rigid editing with Stable Diffusion, aimed at improving the identity preservation quality without compromising editability. Our approach comprises three stages: text optimization, latent inversion, and timestep-aware text injection sampling. Inspired by the success of Imagic, we employ their text optimization for smooth editing. Then, we introduce latent inversion to preserve the input image’s identity without additional model fine-tuning. To fully utilize the input reconstruction ability of latent inversion, we suggest timestep-aware text injection sampling. This effectively retains the structure of the input image by injecting the source text prompt in early sampling steps and then transitioning to the target prompt in subsequent sampling steps. This strategic approach seamlessly harmonizes with text optimization, facilitating complex non-rigid edits to the input without losing the original identity. We demonstrate the effectiveness of our method in terms of identity preservation, editability, and aesthetic quality through extensive experiments.
1 Introduction
Among various image editing tasks, non-rigid editing involves changing the posture or composition of an object while preserving the background and object identity of the original image. It is useful for various applications, such as drawing comic books or videos, where gradual transitions in object actions or compositions over the input image are essential. This task is particularly more challenging than rigid editing, such as global style transfer and image translation, which focus on changing color or textures while retaining the image structure.
The challenges of non-rigid editing become more severe when combined with Stable Diffusion [Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer], primarily related to preserving image identity. Prior works in image editing address this problem through various approaches, including injecting attention maps of the source image into the target image generation process [Hertz et al.(2022)Hertz, Mokady, Tenenbaum, Aberman, Pritch, and Cohen-or, Tumanyan et al.(2023)Tumanyan, Geyer, Bagon, and Dekel, Xiao et al.(2023)Xiao, Yin, Freeman, Durand, and Han, Cao et al.(2023)Cao, Wang, Qi, Shan, Qie, and Zheng, Wang et al.(2024)Wang, Liu, and Xu, Qiao et al.(2024)Qiao, Wang, Su, Zhang, Yu, Wu, and Qi], model finetuning [Ruiz et al.(2023)Ruiz, Li, Jampani, Pritch, Rubinstein, and Aberman, Valevski et al.(2023)Valevski, Kalman, Molad, Segalis, Matias, and Leviathan, Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani], and target text optimization [Gal et al.(2022)Gal, Alaluf, Atzmon, Patashnik, Bermano, Chechik, and Cohen-or, Han et al.(2023)Han, Yang, Kwon, and Ye, Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani]. However, the attention mechanism often leads to unnatural images as it forcefully combines the structure of the source image with the target, resulting in distorted objects and compositions. Also, model finetuning on an input image can lead to a severe overfitting problem and color distortion, especially when applied to Stable Diffusion.
To this end, our work explores improving identity preservation for non-rigid editing with Stable Diffusion by incorporating a text optimization framework. We focus upon the following key insight: text optimization on an input image has shown promise in enabling smooth changes of an object toward the desired editing concept. However, this hinges on the condition that the model faithfully reproduces the input image and ensures that small changes in text embedding do not significantly alter the output image’s resemblance to the input image [Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani].
Therefore, Imagic [Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani] additionally introduced model finetuning, ensuring the model specialized in reproducing the input image. Nevertheless, as this method causes overfitting and color distortion problems, we make a pivotal change by replacing the model fine-tuning with latent inversion. This change maintains the integrity of the original data distribution, preserving the editability of text optimization. Additionally, we exploit the reconstruction power of latent inversion further by inserting source text embedding exclusively during the initial sampling steps. Then, we transition to target text embedding for the subsequent sampling steps, namely timestep-aware text injection. This concept is inspired by the recent study that reveals the pivotal role of the initial sampling steps in determining the coarse structure and outlines of an image. By selectively injecting source and target prompts considering the role of sampling timesteps, we can explore the smooth transition between identity preservation and editability. This approach allows us to find a sweet spot between these two factors.
Our simple yet effective approach effectively harmonizes with text optimization, achieving complex non-rigid editing without losing the original identity. Unlike Imagic, our training-free method offers the advantage of reusing a pre-trained diffusion model regardless of images. Also, our method generates more natural and aesthetically pleasing images than attention-based methods. We compare our edited results both quantitatively and qualitatively with other editing methods, showing superior editing performance on the Stable Diffusion in terms of identity preservation, editability, and aesthetic quality. Moreover, we prove each component of our method is effective through ablation experiments and show the applicability of our method on the different image domain such as Anything-v4.
2 Related Works
2.1 Text-guided Image Manipulation
Text-guided image manipulation involves editing an image based on a given text prompt while preserving the original input image identity. In the realm of this technique, the predominant approach for preserving the original image identity involves the use of an attention mechanism. This mechanism integrates self or cross-attention maps from the input image into the generation process of an edited image [Hertz et al.(2022)Hertz, Mokady, Tenenbaum, Aberman, Pritch, and Cohen-or, Tumanyan et al.(2023)Tumanyan, Geyer, Bagon, and Dekel, Cao et al.(2023)Cao, Wang, Qi, Shan, Qie, and Zheng, Parmar et al.(2023)Parmar, Kumar Singh, Zhang, Li, Lu, and Zhu, Wang et al.(2024)Wang, Liu, and Xu, Qiao et al.(2024)Qiao, Wang, Su, Zhang, Yu, Wu, and Qi]. Additionally, other methods blend noise maps of the source image and target image, thereby maintaining the background from the source image [Couairon et al.(2022)Couairon, Verbeek, Schwenk, and Cord, Avrahami et al.(2022)Avrahami, Lischinski, and Fried, Huang et al.(2023)Huang, Tang, Dong, Lee, and Xu]. Other directions include model finetuning [Ruiz et al.(2023)Ruiz, Li, Jampani, Pritch, Rubinstein, and Aberman, Valevski et al.(2023)Valevski, Kalman, Molad, Segalis, Matias, and Leviathan, Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani, Li et al.(2023)Li, Huang, Ding, and Li] or text optimization [Gal et al.(2022)Gal, Alaluf, Atzmon, Patashnik, Bermano, Chechik, and Cohen-or, Han et al.(2023)Han, Yang, Kwon, and Ye, Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani], which focus on adapting a model or target text to closely match the input image, even to the extent of overfitting. Furthermore, score distillation has been introduced for image editing, where the input image identity is continuously distilled into the target image during the editing process [Hertz et al.(2023)Hertz, Aberman, and Cohen-Or]. Finally, several approaches [Brack et al.(2024)Brack, Friedrich, Hintersdorf, Struppek, Schramowski, and Kersting, Tsaban and Passos(2023), Brack et al.(2023)Brack, Friedrich, Kornmeier, Tsaban, Schramowski, Kersting, and Passos] utilize semantic guidance by adding an edit term in the original classifier-free guidance formula [Ho and Salimans(2021)].
2.2 Diffusion Inversion
For real image editing, the concept of inversion has been widely used to uncover the latent representation that faithfully reconstructs the input real image. DDIM inversion [Song et al.(2020)Song, Meng, and Ermon] excels at reconstructing the original image for unconditional image generation scenarios. However, it faces challenges in conditional cases where the use of classifier-free guidance (CFG) [Ho and Salimans(2021)] magnifies the errors, resulting in decreased editability and reconstruction quality. This drawback of DDIM gave rise to active research in inversion specialized for conditional image generation. Null-text inversion [Mokady et al.(2023)Mokady, Hertz, Aberman, Pritch, and Cohen-Or] improves the inversion quality by optimizing null text embedding $\emptyset$ at each diffusion timestep. On the other hand, Direct Inversion [Ju et al.(2023)Ju, Zeng, Bian, Liu, and Xu] takes a different approach by disentangling source and target diffusion branches, eliminating optimization. This strategy boosts content preservation quality and edit fidelity within a short time. EDICT [Wallace et al.(2023)Wallace, Gokul, and Naik] offers a mathematically exact inversion solution through an alternative inversion process that involves a pair of coupled noise vectors. Edit-Friendly inversion [Huberman-Spiegelglas et al.(2023)Huberman-Spiegelglas, Kulikov, and Michaeli] strongly imprints the input image onto latent noise maps by making them correlated across timesteps, enabling fast and effective editing without the need for optimization.
3 Preliminary
Stable Diffusion Model. Diffusion models aim to estimate a data distribution via a latent variable model. They perform a forward process that gradually adds a Gaussian noise on a clean image at each timestep $t\in[1,...,T]$ until it reaches a noise distribution. Conversely, the reverse process inversely denoises the noise distribution in order to generate a final clean image. The Stable Diffusion model [Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer], an open-source model, conducts these processes on the latent space. Specifically, given an image $\bm{x}\in\mathbb{R}^{H\times W\times 3}$, it encodes $\bm{x}$ into a latent representation $\bm{z}_{0}=\mathcal{E}(\bm{x})$ using a VQ-GAN [Esser et al.(2021)Esser, Rombach, and Ommer] encoder $\mathcal{E}$. The model is trained on this latent space, applying the forward process to $\bm{z}_{0}$. Likewise, $\bm{z}_{0}$ is obtained from the latent variable $\bm{z}_{T}$ using the reverse process. Finally, the generated image is computed by decoding $\bm{z}_{0}$ back to the image space ($\bm{x}=\mathcal{D}(\bm{z}_{0})$). Also, an important function of Stable Diffusion is its capability to facilitate text-guided image generation. This is achieved by inputting a text embedding $\tau_{\phi}(\bm{y})$, which is a text prompt $\bm{y}$ processed by CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.] text encoder $\tau_{\phi}$. The text-guided conditional model is trained using the following objective:
$\small\mathcal{L}_{\text{cond}}=\mathbb{E}_{t,\bm{z}_{0},\bm{\epsilon}}\left[||\bm{\epsilon}-f_{\theta}(\bm{z}_{t},t,\tau_{\phi}(\bm{y}))||^{2}\right],$ (1)
where $f_{\theta}$ is a neural network with parameter $\theta$, $t\sim\texttt{Uniform}[1,T]$, $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$, and $\bm{z}_{t}$ is a noise added version of $\bm{z}_{0}$ using the forward process. Our goal is to improve non-rigid editing with Stable Diffusion, aiming to enhance the accessibility of a publicly available model.
Figure 1: Ablation on Imagic and Our method. The images are edited according to the target text: “A photo of a white horse jumping” and “A photo of a couple hugging on a beach”.
Imagic for Non-rigid Editing. Imagic [Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani], a pioneering method for non-rigid editing, introduced a two-step process: text optimization and model fine-tuning. Firstly, they optimize the target text embedding to best match the input image. Specifically, the target text is encoded to a text embedding $\bm{e}_{tgt}\in\mathbb{R}^{N\times d}$ via a text encoder [Raffel et al.(2020)Raffel, Shazeer, Roberts, Lee, Narang, Matena, Zhou, Li, and Liu], where $N$ is the number of tokens in the target text and $d$ is the token embedding dimension. Following this, they freeze parameters of the pre-trained diffusion model $f_{\theta}$ and optimize the target text embedding $\bm{e}_{tgt}$ using the diffusion model objective:
$\small\bm{e}_{opt}=\arg\min\mathbb{E}_{t,\bm{z}_{0},\bm{\epsilon}}\left[\left\|\bm{\epsilon}-f_{\theta}\left(\bm{z}_{t},t,\bm{e}_{tgt})\right)\right\|^{2}\right].$ (2)
Though the target text embedding is optimized to be matched with the input image, $\bm{e}_{opt}$ often does not show a perfect representation of the input image. Additional optimization does not improve identity preservation but causes a runtime overhead problem (Text optim step 200, Text optim step 1000 in Figure 1). Therefore, Imagic uses model fine-tuning as a firm identity preservation strategy to make the model overfitted to the input image. It freezes the optimized embedding and then fine-tunes the model parameters $\theta$ using the same loss function in Equation 2.
Finally, during the sampling phase, Imagic linearly interpolates the original target text embedding with the optimized one using Equation 3 and passes this through the model to generate an edited image. Here, it can control the degree of editing by varying an interpolation ratio $\alpha$. This aims to establish a smooth transition of text embedding toward the desired editing concept while maintaining the original identity.
$\bm{e}_{int}=\alpha\cdot\bm{e}_{tgt}+(1-\alpha)\cdot\bm{e}_{opt}$ (3)
It is worth noting that Imagic is particularly effective when applied to Imagen [Saharia et al.(2022)Saharia, Chan, Saxena, Li, Whang, Denton, Ghasemipour, Gontijo Lopes, Karagol Ayan, Salimans, et al.], a large-scale generative model only accessible for business. It successfully applies the editing concept while faithfully maintaining the original identity and background of the original image. However, the same recipe encounters limitations when applied to Stable Diffusion, as mentioned in previous works [Han et al.(2023)Han, Yang, Kwon, and Ye, Brack et al.(2023)Brack, Friedrich, Kornmeier, Tsaban, Schramowski, Kersting, and Passos]. Our experiments also affirm the unsatisfactory editing performance of Imagic on Stable Diffusion (Sec 5), with a noticeable deficiency in identity preservation.
Figure 2: Overview of our method.
4 Method
In this section, we identify the problem of Imagic combined with Stable Diffusion and introduce our key contributions to enhancing non-rigid editing performance. An overview of our method is demonstrated in Figure 2.
We follow the text optimization from Imagic to preserve the original identity and achieve smooth non-rigid editing (See Figure 2 (A)). As text optimization alone is insufficient for a perfect match with the input image, Imagic solves this problem via model finetuning. However, in the Stable Diffusion setting, we found that the model finetuning presents several significant drawbacks that adversely affect editing quality. Firstly, it triggers a severe tradeoff between (1) overfitting, where the model overly memorizes the original image and thus cannot faithfully incorporate the text prompt well, or (2) the lack of identity preservation, where the edited image aligns with the target text but it fails to preserve the original image’s identity. Figure 1 provides visual evidence that the edited image closely resembles the input image as we finetune the model longer. Conversely, minimal finetuning produces an arbitrary image that bears little resemblance to the input image, indicating the lack of identity preservation (Text optim step 200 + Finetune step 100, +Finetune step 400 in Figure 1). Model finetuning thus hardly satisfies both identity preservation and editability simultaneously. (More ablation experiments on finetuning steps are in Appendix.) Even worse, it causes color distortion, disrupting the data distribution. This is because model training aims to reach the data distribution of the training dataset, which in this context consists of only a single input image.
Instead of model fine-tuning, we utilize latent inversion to maintain the original identity (See Figure 2 (B)). The inversion process is defined by Equation 4:
$\small\bm{z}^{*}_{t+1}=$ $\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}\bm{z}^{*}_{t}+$
$\sqrt{\alpha_{t+1}}\left(\sqrt{\frac{1}{\alpha_{t+1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\right)f_{\theta}(\bm{z}^{*}_{t},t,\bm{e}_{src}),$ (4)
where $\alpha_{t}$ is a noise schedule at timestep $t$ and $\bm{e}_{src}$ is a source text embedding. As the inversion produces a latent that faithfully reconstructs the original image, the use of inverted latent for sampling can effectively maintain the original identity with text-guided editing. Moreover, the inversion method no longer suffers from (1) the tradeoff between overfitting and editability and (2) color distortion.
While employing the inverted noise during sampling process helps maintain the overall color and context of the original image, introducing a new text prompt different from the source text can cause deviations in image configuration, such as the size and position of the object, compared to the original image. To address this issue, we leverage a well-known observation that diffusion models generate global structures at the initial timesteps and fine details toward the end [Choi et al.(2022)Choi, Lee, Shin, Kim, Kim, and Yoon]. Specifically, we introduce an additional regularization depending on sampling timesteps (See Figure 2 (C)). During the initial steps ($n\rightarrow T$), we incorporate the source text embedding and then transition to interpolated embeddings for the subsequent timesteps:
[…]
This strategy helps to closely adhere to the global structure of the original image, a key component of the image identity, during the early stage of the editing process. Since then, the editing concept is applied to the subsequent sampling steps, effectively incorporating editing concepts without altering the image identity. We show its substantial impact on preserving the identity through ablation studies in Section 5.4.
5 Experiments
5.1 Implementation Details
Baselines. For a comprehensive comparison, we evaluate our method against the following competitive techniques: Imagic [Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani], Null-text inversion [Mokady et al.(2023)Mokady, Hertz, Aberman, Pritch, and Cohen-Or], DDS [Hertz et al.(2023)Hertz, Aberman, and Cohen-Or], Masactrl [Cao et al.(2023)Cao, Wang, Qi, Shan, Qie, and Zheng] and Edit-Friendly [Huberman-Spiegelglas et al.(2023)Huberman-Spiegelglas, Kulikov, and Michaeli]. Notably, Imagic and Masactrl are specialized for the non-rigid editing task. Null-text inversion and Edit-Friendly represent prominent methods designed for natural image editing via inversion. DDS proposes a new paradigm of image editing using score distillation. In our evaluations, we run the released codes of all baseline models using pre-trained Stable Diffusion v1.4 with publicly available checkpoints. Unfortunately, we could not include other non-rigid editing methods, such as LayerDiffusion [Li et al.(2023)Li, Huang, Ding, and Li] and BARET [Qiao et al.(2024)Qiao, Wang, Su, Zhang, Yu, Wu, and Qi], as their codes are not released as of our submission date. Details on experiments are in Appendix.
Evaluation Metrics. We utilize CLIP-T score [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.] for measuring target text alignment and LPIPS [Zhang et al.(2018)Zhang, Isola, Efros, Shechtman, and Wang], PSNR, and MSE for measuring input image identity preservation quality of edited images. However, these metrics are incapable of evaluating the aesthetic quality of the edited image; they cannot penalize the distorted or unnatural outputs. To mimic the standard of human visual inspection, we additionally employ Aesthetic score [Schuhmann(2022)], Pickscore [Kirstain et al.(2024)Kirstain, Polyak, Singer, Matiana, Penna, and Levy], and Imagereward [Xu et al.(2024)Xu, Liu, Wu, Tong, Li, Ding, Tang, and Dong] which measure how much people like the image on average.
Figure 3: Qualitative comparison with state-of-the-art methods on Stable Diffusion v1.4.
5.2 Qualitative Evaluation
We compare the editing results of our method to state-of-the-art methods using a pre-trained Stable Diffusion v1.4 in Figure 3. We show our results using Null-text inversion as it provides the best editing performance (Figure 5). For a fair comparison, we also applied Null-text inversion in Masactrl, which originally used DDIM inversion. It results in improved identity preservation and editability of Masactrl. We tried various hyperparameters of each method and chose the best result for each.
Firstly, Null-text inversion alone for editing, in which we just input an inverted latent and a target text through the model, shows low identity preservation. It does not retain the original structure of the input image and simply pursues generating an image that aligns well with the text prompt. This is expected because there is no other constraint for identity preservation during the sampling process. Meanwhile, Imagic has a serious tradeoff between overfitting and editability. Model finetuning especially causes a severe overfitting problem, which distorts the data distribution, whereas a random noise without finetuning during the sampling process ruins the structure preservation. Despite the extensive hyperparameter search, it is hard to find a sweet spot between these two factors (See Appendix). Besides, DDS also suffers from overfitting to the input image, especially when handling non-rigid editing. This shows that score distillation has trouble changing geometry. On the other hand, Edit-Friendly and Masactrl attempt to apply an editing concept while maintaining the original identity by improving latent noise space and utilizing attention maps, respectively. When an image has no complicated background and contains a single object, such as a Spiderman image in the third row of Figure 3, their editing quality is sufficiently high. However, they show unnatural edited results in other cases. In the first row of Figure 3, a bulldog opens its mouth but does not smile in Edit-Friendly, and the bulldog shows a grotesque mouth in Masactrl. In the second row, a horse jumps in both methods, but the hind legs of the horse remain similar to the input image, resulting in an unnatural jumping appearance. Overall, our method clearly shows the editing concept while faithfully preserving the original identity, like a bulldog with a big smile and a horse jumping dynamically. These visual trends and the superiority of our method consistently hold in other results. More qualitative results are demonstrated in Appendix, including results using the pretrained Anything-v4 model.
5.3 Quantitative Evaluation
Due to inherent inaccuracies and ambiguities in quantitative evaluation of non-rigid editing, we conduct a user study. We utilized the TedBench dataset [Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani] to generate 50 edited images of each baseline model. A set of input real image, target text and 6 edited images using each model is shown to the user. The user is asked to choose one among edited images that best satisfies both two conditions simultaneously: The edited image 1) maintains background and appearance of the object in the original input image and 2) aligns with the target text. We randomly shuffled the order of edited images so that the user doesn’t know a model used to generate each image. We accumulated 38 respondents, resulting in total 1900 (38 $\times$ 50) responses. Figure 4 (A) shows the preference rate of each method, which represents the proportion of times the corresponding method was selected out of 1900 responses. Ours receives the highest preference rate, 0.43, which is more than twice as high as the preference rate of the second-ranked Masactrl (0.19). It indicates that our method surpasses other models in terms of both similarity to the input image and text alignment.
We further conduct a quantitative evaluation on the models that ranked within the top three in user study: Edit-Friendly, Masactrl, and Ours. We utilized the same target text prompts used in user study to generate 50 edited images, producing five variants for each edited image with different seeds, resulting in a total of 250 images. In Figure 4 (B), we present various metrics of each method at different CLIP-T scores, where CLIP-T changes depending on hyperparameters. Generally, Edit-Friendly receives the best score in LPIPS ($\downarrow$), PSNR ($\uparrow$) and MSE ($\downarrow$) showing that it has a strong tendency to preserve the identity of the original image. Compared to Edit-Friendly, Ours and Masactrl edit images in a way that align more closely with the text prompt. For non-rigid editing which involves changes in motion of objects, big changes in the structure of objects should be allowed. Therefore, strictly preserving the structure of the original image disturbs editing. Our user study proves that our method well balances input image similarity and editability. This success can be attributed to our inversion and timestep-aware text injection sampling, which preserve the original identity without disrupting the data distribution. This enables text optimization to smoothly transition towards the desired editing concept while maintaining the background and object identity faithfully. Also, our method receives the highest score in Aesthetic score ($\uparrow$) and Imagereward ($\uparrow$) indicating that our edited images are more preferred in terms of human preference. Masactrl, on the other hand, presents the lowest score in Aesthetic score, Pickscore and Imagereward in general, consistent with its tendency to generate unnatural images with distorted object appearances and poses. This can be attributed to the difference in object appearance and poses between the source and target images, each generated using the original source and target text, respectively. Example images are available in Appendix. As the appearance gap between the target image and the source image intensifies, inserting the attention maps of the source image into the target image sampling process often results in abnormal objects with unnatural appearances and poses.
Figure 4: User study and quantitative comparison.
5.4 Ablation Experiments
Inversion Type. Figure 5 demonstrates the impact of different inversion methods: DDIM [Song et al.(2020)Song, Meng, and Ermon] and Null-text inversion [Mokady et al.(2023)Mokady, Hertz, Aberman, Pritch, and Cohen-Or]. Each image is edited with the target text “A photo of a jumping dog.” When employing DDIM inversion for our method, it is highly close to the input image but has low editability. This again proves that DDIM inversion is suboptimal for real image editing applications. On the other hand, Null-text inversion is known to improve reconstruction and editing quality, particularly when using classifier-free guidance with a guidance scale greater than 1, as compared to DDIM inversion. Interestingly, when inputting a source prompt (e.g. “A photo of a dog") following the original method, the preservation quality sometimes significantly deteriorates while focusing on editing (NTI - src). Conversely, when we set a source text prompt equal to null text (‘’), the edited images show significant improvement in structural preservation (NTI - $\emptyset$). Therefore, when the reconstruction performance falls behind, employing the null text for the source prompt can enhance the overall reconstruction ability. Overall, Null-text inversion demonstrates superior performance in editing quality to DDIM.
Figure 5: Ablation on inversion type.
Source Text Injection Ratio. Our timestep-aware sampling strategy injects the source text embedding at the beginning of the sampling process instead of an interpolated embedding to improve identity preservation. Figure 1 includes the ablation study with different source text input ratios. Given that we use DDIM sampling with 50 steps, inserting the source text for the initial 10 steps corresponds to a 20% source text input ratio. We present results under three different scenarios: a ratio of 0%, 10%, and 20%, depicted as Text optim step 200 + Inversion, + Src ratio 10%, and + Src ratio 20% in Figure 1, respectively. Notably, as the source text input ratio increases, the edited images increasingly retain the original structure of the input image. We empirically observe that inserting the source text about 5 to 20 steps, which corresponds to a ratio of 10% to 40%, leads to the most favorable editing results depending on the degree of desired structural changes.
6 Conclusion
Our work has advanced the capabilities of non-rigid editing with Stable Diffusion, particularly focusing on identity preservation. Our method comprises text optimization, latent inversion, and timestep-aware text injection sampling. Based on the analysis of the previous study, we chose text optimization as the baseline as it allows a smooth transition toward a desirable editing concept. To overcome the limitations of model fine-tuning from Imagic, we suggest utilizing the reconstruction power of latent inversion. To further improve the identity preservation quality, we introduce timestep-aware text injection sampling that selectively applies source and target prompts depending on sampling timesteps. Our method presents superior performance in both qualitative and quantitative comparisons with other editing methods on Stable Diffusion. Our method does not trigger overfitting tradeoffs and color distortion, unlike model finetuning in Imagic. Also, we can provide natural and aesthetically pleasing images compared to attention-based methods. Finally, we can extend the usage of our method for other image domains like Anything-V4.